# The Basis of Foot-Sortable Sock Orderings

### v1

@@ 4. Prelimanary Observations and Lemmas
Note that, by the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$. Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$. Therefore, Lemma 24 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$. Note that this means that $z=t_{n}$ since $t_{n}$ is defined to be the last color that appears directly above $d$ on the stack when sorting $X$ with $\theta_{z}$. Therefore,
$X\supseteq aydI(a,t_{0},\dotsc,t_{n})R_{2}\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})yz=aydt_{0}aI(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
@@
The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$. Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 6. In Lemma 27 and 28, we only need the first conclusion “$X$ is unsortable”, to prove that the algorithm is correct. The second conclusion “a subpattern of $X$ is in $\Gamma$” is needed in our inductive proof of Theorem 1.
@@
  • $atb$ or $abt$ is contained in $R$ and $a$ is blocked by a $d$-sandwich in $X$; or
@@
then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
@@
Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$, where $d$ is the last color sorted by the sequence. Further suppose that the following two conditions hold:
@@
Then there exists $P\subseteq R$ such that $a$ is the only good sortable color in $atP$. Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable. If, in addition, every strict subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$. Furthermore, two of the three colors can be the top two colors on $S$ (cf. Algorithm 3, Line 8)
@@
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and $S$ contains at least three colors. Suppose $(S,R)$ is terminal. If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
@@
Suppose $(S,R)$ is a counterexample. We can assume without loss of generality that no colors are repeated in $S$. So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors. Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal. Since $(S,R)$ is a counterexample, we have $p\geq 2$, and if we let $W$ be the nonempty string $y_{3}\dotsm y_{p}b$, then for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$ and $y_{i}z_{i}\nsubseteq R$. Since $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$, Lemma 24 implies that $z_{1}$ is unsortable in $(y_{2}Wt,R)$. Therefore, because $y_{1}z_{1}\nsubseteq R$ and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$. If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ that does not contain $y_{1}$, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. So, in both cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$, so $y_{1}\neq z_{2}$, and Lemma 24 then implies that $y_{1}z_{2}\subseteq R$. Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$. Therefore, $y_{2}\neq z_{1}$. This together with Lemma 24 implies that $y_{2}z_{1}\subseteq R$. Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$. But we have already established that $y_{1}z_{1}\nsubseteq R$. ∎
Our last lemma is simple despite the somewhat technical statement. It is used in the case when a color, say $x$, is not sorted by the greedy algorithm despite terminating before the color which is sorted by the greedy algorithm, say $y$. Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable. The lemma deals with the other possibility. Namely, $x$ already being beneath the top color on the stack.
@@
Let $S$ and $R$ be sock orderings where $S$ does not contain a sandwich and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$. Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable. If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ precedes $x$ on $R$.
@@
Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich. Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$. The fact that $x$ is unsortable then implies that $x$ is below the top element on the stack $S$. Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$. Therefore, the colors $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ can precede $x$ on $R$. ∎
5. Proof of Theorem 1
The proof is by induction on $|X|$. The base case is vacuously true. So, by the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable. Note that if we have colors $a$, $x$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 25, then we can assume (I) holds. That is, we can assume that $a$ is blocked by a $d$-sandwich. This is because (II) and (III) in Lemma 25 cannot hold, and if (IV) holds, then we satisfy the conclusion. This will make it easy to apply either Lemma 27 or 28 to complete the proof.
Let $(S,R):=\phi_{G}(X)$ be the terminal state of the greedy algorithm and let $d$ be the last color sorted by $\phi_{G}$. Since $X$ is unsortable, we have $R\neq\emptyset$.
Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,S)$ is unsortable. If $S=\emptyset$, then Lemma 26 implies that a subpattern of $R\subseteq X$ is in $\Gamma$. Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable, and Lemma 26 implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
From now on, we will assume that we are not in Case 1. That is, for every $x\in S$, some color in $xR$ is sortable. Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$. Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$.
Case 2: $(bt,R)$ is terminal. By Lemma 28, there exists $P\subseteq R$ such that $b$ is the only good sortable color in $btP$ and that if $b$ is blocked by a $d$-sandwich in $X$, then a subpattern of $X$ is in $\Gamma$. If $b$ is not blocked by a $d$-sandwich in $X$, then we must satisfy Lemma 25(IV). That is, $I_{1,n}\preceq X$, for some $n\geq 0$.
Case 3: $(bt,R)$ is not terminal. With Lemma 29, we can define $a$ to be the lowest color in $S$ such that $(abt,R)$ is terminal.
@@ 5. Proof of Theorem 1
Suppose, by way of contradiction, that $at\nsubseteq R$. Then, because $t$ is unsortable in $(abt,R)$, Lemma 24 implies that $t$ is unsortable in $(bt,R)$. Since $(bt,R)$ is not terminal and both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$. If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$. If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 24 implies that $ay\subseteq R$. Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$. So, in both cases, $y$ follows the last occurrence of $t$ in $btR$. This with Lemma 22 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction. ∎
Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$. By Claim 31, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$. But then Lemma 25 and Lemma 27 lead to a contradiction, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$. Recall that $abt\subseteq S$. This and the case, implies that $b\notin R$. Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$. Because $b\notin R$ there exists a color in $C$ that is sorted during $\phi_{G}(X)$. Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted by $\phi_{G}(X)$. Note that we have
@@
Furthermore, by the definition of the greedy algorithm, the color $b$ can never appear as the top element of the stack after $c$ is sorted by the greedy algorithm. For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm. Note that then $t=t_{n}$, so Claim 31 implies that $at_{n}=at\subseteq R$. With Lemma 21, we have
@@
There are at least three colors in $S^{\prime}$. More specifically, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
@@
First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Then, because the greedy algorithm sorted $c$ and not $b$ at stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable. Then, Lemma 30 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$. Since (1), (2), and (4), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$. By (3), we also have
@@
Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$; we will show that $\hat{y}=y$ and our conclusion holds. To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$. Furthermore, (2) and (4) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
@@
If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (5), we will also have $a\hat{y}b$ on the stack. Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm. Furthermore, since we cannot have $\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$. But this contradicts the fact that we are in Case 3.2.
Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$. First note that since $c$ is sortable with respect to $(S^{\prime},A^{\prime}bC)$, the color $c$ is not below the top color on the stack $S^{\prime}$. Since there is a sandwich consisting of $\hat{y}$ and $c$ in $S^{\prime}AbC$, we must have $\hat{y}$ in $A^{\prime}bC$. Also, by the definition of $c$, we have $c\in C\subseteq A^{\prime}bC$. Since the greedy algorithm sorts $c$ and not $\hat{y}$ when $(S^{\prime},A^{\prime}bC)$ appears, $c$ is sortable and $\hat{y}$ is unsortable with respect to $(S^{\prime},A^{\prime}bC)$. Furthermore, since the selection of $\hat{y}$ implies that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$, Lemma 30 implies that $\hat{y}$ and $c$ are the top two colors in $S^{\prime}$. Note that we have shown that $\hat{y}=y$. Furthermore, Lemma 30 implies that and no color except possibly $c$ precedes $\hat{y}=y$ in $A^{\prime}bC$. In particular, $b$ does not precede $\hat{y}=y$ in $A^{\prime}bC$, so $\hat{y}=y\in A^{\prime}$. Then, with (3),

### v2

@@ 2.4. Observations and Lemmas
Note that since $y$ is immediately below $d$ on the stack after the sequence $\theta_{a}$ is applied to $X$ and $d$ is never removed from the stack while $\theta_{z}$ is applied to $X$, the color $y$ is immediately below $d$ on the stack $S_{2}$. Let $S_{2}^{\prime}:=S_{2}-d$ and note that $(S_{2}^{\prime},R_{2})=\theta_{z}(X-d)$. By the selection of $z$, the color $z$ is unsortable in $(S_{2},R_{2})$, but $z$ is sortable in $(S^{\prime}_{2},R_{2})$. Since $d\notin S^{\prime}_{2}R_{2}\subseteq X-d$, we have $z\neq d$ and $dz\nsubseteq R_{2}$, so Lemma 7 implies that $y$, $d$, and $z$ are the top three colors in the stack $S_{2}$ and that $yz\subseteq R_{2}$. Note that this means that $z=t_{n}$, since $t_{n}$ is defined to be the last color that appears immediately above $d$ on the stack when we sort $X$ with $\theta_{z}$. So, $yt_{n}\subseteq R_{2}$. This, with (1) and (2), implies that
$X=YdW\supseteq aydt_{0}aI(t_{0},\dotsc,t_{n})R_{2}\supseteq aydIt_{0}a(t_{0},\dotsc,t_{n})yt_{n}\cong I_{1,n}.\qed$
@@
The following three lemmas are needed in the proof of our main theorem to show that a given unsortable sock ordering has a subpattern that is in $\Gamma$. Unsurprisingly, these proofs have a number of cases and are somewhat technical, so we defer their proofs until Section 2.6.
@@
  • $atb\subseteq R$ or $abt\subseteq R$ and $a$ is blocked by a $d$-sandwich in $X$; or
@@
then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
@@
Let $X$ be a sock ordering and suppose that a nonempty sorting sequence produces $(S,R)$ when it is applied to $X$ where $d$ is the last color sorted by the sequence. Further suppose that the following two conditions hold:
@@
Then there exists $P\subseteq R$ such that $a$ appears in $P$ and $a$ is the only good sortable color in $atP$. Furthermore, if $a$ is blocked by a $d$-sandwich in $X$, then $X$ is unsortable. If, in addition, every proper subordering of $X$ is sortable, then a subpattern of $X$ is in $\Gamma$.
The following lemma shows that if we reach the state $(S,R)$ that is terminal and $S$ has at least three colors, then, in a certain sense, we only need to focus on at most three colors in $S$. Furthermore, two of the three colors can be the top two colors on $S$.
@@
Let $S$ and $R$ be sock orderings where $S$ is sorted and at least three distinct colors appear in $S$. Suppose $(S,R)$ is terminal. If $b$ and $t$ are the top two colors on $S$, then there exists a color $a\in S$ distinct from $b$ and $t$ such that $(abt,S)$ is terminal.
@@
Suppose $(S,R)$ is a counterexample. We can assume without loss of generality that no colors are repeated in $S$. So we can write $S=x_{1}x_{2}\dotsm x_{m}bt$ where $x_{1},\dotsc,x_{m}$, $b$, and $t$ are distinct colors. Let $p$ be as small as possible such that there exists $y_{1},\dotsc,y_{p}$, a subsequence of $x_{1},\dotsc,x_{m}$, such that $(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal. Since $(S,R)$ is a counterexample, we have $p\geq 2$. If $p=2$, let $W=b$ and otherwise let $W=y_{3}\dotsm y_{p}b$. By the minimality of $p$, for $i\in\{1,2\}$, there exists some $z_{i}$ that is sortable in $(y_{i}Wt,R)$.
For $i\in\{1,2\}$, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, we have $z_{i}\notin y_{i}W$. In particular, this implies that $z_{i}\neq y_{i}$ and $z_{i}\neq b$. Furthermore, since $z_{i}$ is sortable in $(y_{i}Wt,R)$, there is no $y_{i}by_{i}$ sandwich blocking $z_{i}$ in $y_{i}WtR$, so $y_{i}z_{i}\nsubseteq R$.
We will now show that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. Recall that $(y_{1}y_{2}Wt,R)=(y_{1}y_{2}\dotsm y_{p}bt,R)$ is terminal, so $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$. Therefore, we can apply Lemma 7 to deduce that $z_{1}$ is unsortable in $(y_{2}Wt,R)$, because $z_{1}\neq y_{1}$, $y_{1}z_{1}\nsubseteq R$, and $y_{2}$ is not the second color on the stack $y_{1}y_{2}Wt$. Since $z_{1}$ is unsortable in $(y_{2}Wt,R)$, $y_{1}z_{1}\nsubseteq R$, and $z_{1}\notin W$, either $z_{1}=y_{2}$ or there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear. If $z_{1}=y_{2}$, then, because $z_{1}=y_{2}$ is sortable in $(y_{1}Wt,R)$, we have $y_{2}\in R$, so $y_{1}y_{2}=y_{1}z_{1}\nsubseteq R$ implies that either $y_{1}\notin R$ or that $y_{1}$ is blocked by a sandwich of the form $y_{2}by_{2}$ in $y_{2}WtR$. So, $y_{1}$ is not sortable in $(y_{2}Wt,R)$. If there is a sandwich that blocks $z_{1}$ in $y_{2}WtR$ in which $y_{1}$ does not appear, then $y_{1}z_{1}\nsubseteq R$ implies that $y_{1}$ is not sortable in $(y_{2}Wt,R)$. So, in all cases, $y_{1}$ is not sortable in $(y_{2}Wt,R)$.
Since $y_{1}$ is not sortable in $(y_{2}Wt,R)$, we have $y_{1}\neq z_{2}$. This, with the fact that $z_{2}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{2}Wt,R)$ together with Lemma 7, yields $y_{1}z_{2}\subseteq R$. Since $y_{1}z_{2}\subseteq R$ and $y_{2}z_{2}\nsubseteq R$, $y_{2}$ is not sortable in $(y_{1}Wt,R)$, so $y_{2}\neq z_{1}$. Using this fact with Lemma 7, allows us to conclude that $y_{2}z_{1}\subseteq R$ because $z_{1}$ is not sortable in $(y_{1}y_{2}Wt,R)$ but is sortable $(y_{1}Wt,R)$. Since $y_{1}z_{2}\subseteq R$, $y_{2}z_{2}\nsubseteq R$, and $y_{2}z_{1}\subseteq R$, we have $y_{1}z_{1}\subseteq R$. But we have already established that $y_{1}z_{1}\nsubseteq R$. ∎
Our last lemma is simple despite the somewhat technical statement. Recall the greedy algorithm described in Section 2.3. This lemma is used in the case when a color, say $x$, is not sorted by the greedy algorithm at a particular stage despite terminating before another color, say $y$, which is sorted by the greedy algorithm at that stage. Note that, in this case, if a sandwich blocks $x$, then it must be a $y$-sandwich because $y$ is sortable. The lemma deals with the other possibility. Namely, $x$ already being beneath the top color on the stack.
@@
Let $S$ and $R$ be sock orderings where $S$ is sorted and let $x$ and $y$ be distinct colors in $R$ such that the final $x$ in $R$ appears before the final $y$ in $R$. Suppose that, with respect to $(S,R)$, the color $y$ is sortable and the color $x$ is unsortable. If no $y$-sandwich blocks $x$ in $SR$, then $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ precedes $x$ in $R$.
@@
Because $y$ is sortable in $SR$, every sandwich in $SR$ that appears before the final $y$ in $R$ must be a $y$-sandwich. Therefore, because the final $x$ in $R$ precedes the final $y$ in $R$ and $x$ is not blocked by a $y$-sandwich in $SR$, the color $x$ is not blocked by a sandwich in $SR$. The fact that $x$ is unsortable then implies that $x$ is below the top color on the stack $S$. Since $y$ is sortable in $SR$, there is no $x$-sandwich that blocks $y$ in $SR$. Therefore, the colors $x$ and $y$ are the top two colors on $S$. Furthermore, no color except possibly $y$ can precede $x$ in $R$. ∎
2.5. Proof of Theorem 3
The proof is by induction on the number of socks in $X$. The base case is vacuously true. So assume that $X$ is an unsortable sock ordering. With the induction hypothesis, we can assume that there does not exist a proper subordering of $X$ that is unsortable.
Note that if we have colors $a$ and $d$ and substrings $S^{\prime}\subseteq S\subseteq X$ and $R^{\prime}\subseteq R\subseteq X$ that satisfy the hypothesis of Lemma 8, then we can assume Lemma 8(I) holds. That is, we can assume that $a$ is blocked by a $d$-sandwich. To see this, note first that Lemma 8(II) cannot hold because every proper subordering of $X$ is sortable. Additionally, Lemma 8(III) cannot hold because $X$ is unsortable. Furthermore, if Lemma 8(IV) holds, then we satisfy the conclusion. This will make it possible to apply either Lemma 10 or 11 to complete the proof.
Let $(S,R)$ be output of the greedy algorithm (as described in Section 2.3) applied to $X$. This implies that $(S,R)$ is terminal. Since $X$ is unsortable, we have $R\neq\emptyset$.
Case 1: $S=\emptyset$ or there exists $x\in S$ such that $(x,R)$ is terminal. If $S=\emptyset$, then $R$ is trivially unsortable, so Lemma 9 implies that a subpattern of $R\subseteq X$ is in $\Gamma$. Similarly, if there exists $x\in S$ such that $(x,R)$ is terminal, then $xR$ is trivially unsortable. Lemma 9 then implies that a subpattern of $xR\subseteq X$ is in $\Gamma$.
From now on, we will assume that we are not in Case 1. That is, for every $x\in S$, some color in $xR$ is sortable. Since $(S,R)$ is terminal, this implies that there are at least two colors on $S$. Define $t$ to be the top color in $S$, and let $b$ be the color in $S$ appearing immediately under $t$. We can also now let $d$ be the last color sorted by the greedy algorithm.
Case 2: $(bt,R)$ is terminal. By Lemma 11, there exists $P\subseteq R$ such that $b$ appears in $P$ and such that $b$ is the only good sortable color in $btP$. This means that the hypothesis of Lemma 8 is satisfied with $b$, $d$, $bt$, and $P$ playing the roles of $a$, $d$, $S^{\prime}$, and $R^{\prime}$, respectively. Therefore, we can assume that $b$ is blocked by a $d$-sandwich in $X$. But then Lemma 11 implies that a subpattern of $X$ is in $\Gamma$.
Case 3: $(bt,R)$ is not terminal. With Lemma 12, we can define $a$ to be the lowest color on the stack $S$ such that $(abt,R)$ is terminal.
@@ 2.5. Proof of Theorem 3
Suppose, by way of contradiction, that $at\nsubseteq R$. Then, because $t$ is unsortable in $(abt,R)$ and $t\neq a$ and $a$ is not the second color on the stack $abt$, Lemma 7 implies that $t$ is unsortable in $(bt,R)$. By the case, $(bt,R)$ is not terminal, so since both $b$ and $t$ are not sortable in $(bt,R)$, there must exist a color $y\in R$ that is not $b$ or $t$ which is sortable in $(bt,R)$. If $y=a$, then, because $at\nsubseteq R$, the color $y=a$ must follow the last occurrence of $t$ in $btR$. If $y\neq a$, then, since $y$ is sortable in $(bt,R)$ and unsortable in $(abt,R)$, Lemma 7 implies that $ay\subseteq R$. Because $at\nsubseteq R$, this implies that $y$ follows the last occurrence of $t$ in $btR$. So, in both cases, $y$ follows the last occurrence of $t$ in $btR$. This with Lemma 5 and the fact that $t$ is the second color in $btR$ and $t$ is unsortable in $btR$ give us that $y$ is unsortable in $btR$, a contradiction. ∎
Case 3.1: There exists a color $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$ and $z\in R$. By Claim 14, we can assume that either $atz$, $azt$, or $zat=zaI(t)t$ is in $R$. But then Lemma 8 and Lemma 10 give us the desired conclusion, because $a$ is the only good sortable color in both $aztatz$ and $aztazt$, and $z$ is the only good sortable color in $aztzat$.
Case 3.2: For every $z\in S$ distinct from $a$ and $t$ such that $azt\subseteq S$, we have $z\notin R$. Recall that $abt\subseteq S$. This and the case, implies that $b\notin R$. Let $A$ and $C$ be such that $X=AbC$ and $b\notin C$. Because $b\in S$ and $b\notin R$, there exists a color in $C$ that is sorted by the greedy algorithm. Let $c\in C$ be the color in $C$ that is sorted first by the greedy algorithm and let $(S^{\prime},A^{\prime}bC)$ be the state immediately before $c$ is sorted. Note that we have
@@
Furthermore, since $b$ is not sorted by the greedy algorithm, the color $b$ can never appear as the top color on the stack after $c$ is sorted by the greedy algorithm. For $n\geq 0$, let $t_{0},\dotsc,t_{n}$ be the $n+1$ distinct colors that appear immediately above $b$ after $c$ is sorted by the greedy algorithm. Note that then $t=t_{n}$, so Claim 14 implies that $at_{n}=at\subseteq R$. With Lemma 4, we have
@@
There are at least three colors in $S^{\prime}$. Furthermore, if $y$ is the color immediately beneath the top color in $S^{\prime}$, we have $ayc\subseteq S^{\prime}$ and $yI(c,t_{0},\dotsc,t_{n})at_{n}\subseteq A^{\prime}bC$.
@@
First assume that $b$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Then, because the greedy algorithm sorted $c$ and not $b$ at the stage $(S^{\prime},A^{\prime}bC)$ and the final $b$ occurs before the final $c$ in $S^{\prime}A^{\prime}bC$, with respect to $(S^{\prime},A^{\prime}bC)$, the color $b$ is unsortable and the color $c$ is sortable. Then, Lemma 13 implies that $b$ and $c$ are the top two colors on the stack $S^{\prime}$, so $y=b$. Since (3), (4), and (6), all hold and $b$ and $c$ are the top two colors on the stack $S^{\prime}$, we must have $ayc=abc\subseteq S^{\prime}$. By (5), we also have
@@
Now assume that $b$ is blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Let $\hat{y}$ be the other color in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$. We will show that $\hat{y}=y$ and our conclusion holds. To see this, first note that we must have $\hat{y}c\subseteq S^{\prime}A^{\prime}\subseteq A$ and $c\hat{y}\subseteq S^{\prime}A^{\prime}\subseteq A$. Furthermore, (4) and (6) and $c\hat{y}\subseteq S^{\prime}A^{\prime}$ imply that $\hat{y}\neq a$ and
@@
If the final $\hat{y}$ follows the final $c$ in $S^{\prime}A^{\prime}bC$, then, after $c$ is sorted, the color $\hat{y}$ must appear off the stack and, with (7), we will also have $a\hat{y}b$ on the stack. Then, the fact that $b$ is never sorted by the greedy algorithm implies that $\hat{y}$ is never sorted by the greedy algorithm. Furthermore, since we cannot have a $\hat{y}b\hat{y}$-sandwich on the stack, we must have $a\hat{y}bt_{n}\subseteq S$ and $\hat{y}\in R$. But this contradicts the fact that we are in Case 3.2.
Therefore, the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$. Recall that $\hat{y}$ is the color other than $c$ that is in the first $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$. Since $b\notin C$, we have $\hat{y}b\subseteq S^{\prime}A^{\prime}b$. Because there is no sandwich consisting of $b$ and $\hat{y}$ which blocks $c$ in $S^{\prime}A^{\prime}bC$ and both the final $b$ and the final $\hat{y}$ in $S^{\prime}A^{\prime}bC$ preceed the final $c$ in $S^{\prime}A^{\prime}bC$, this implies that $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$. In particular, we have $\hat{y}\notin C$. We claim that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. For a contradiction, assume that $\bar{y}$ is the other color in such a sandwich. Since $\hat{y}\notin C$, this sandwich is in $S^{\prime}A^{\prime}b$, and $\bar{y}\hat{y}\subseteq S^{\prime}A^{\prime}b$. Since $b\hat{y}\nsubseteq S^{\prime}A^{\prime}bC$, this implies that $\bar{y}\neq b$. Therefore, there is a sandwich consisting of $\bar{y}\neq b$ and $c$ in $S^{\prime}A^{\prime}b$, so this sandwich blocks $b$ as well as $\hat{y}$. Furthermore, since there is no sandwich consisting of $\bar{y}$ and $\hat{y}$ that blocks $c$ in $S^{\prime}AbC$, we have $\hat{y}\bar{y}\nsubseteq S^{\prime}Ab$. But this contradicts the fact that $\hat{y}$ is in the first $c$-sandwich in $S^{\prime}A^{\prime}bC$ that blocks $b$.
So we have established that the color $\hat{y}$ is not blocked by a $c$-sandwich in $S^{\prime}A^{\prime}bC$. Recall that the final $\hat{y}$ precedes the final $c$ in $S^{\prime}A^{\prime}bC$, so Lemma 13 implies that $c$ is on top of the stack $S^{\prime}$ and the color $\hat{y}$ is immediately below the color $c$ on $S^{\prime}$. Note that we now have $\hat{y}=y$ since $y$ was defined to be the color below the top color on $S^{\prime}$. Since $y=\hat{y}$ is in a $c$-sandwich that blocks $b$ in $S^{\prime}A^{\prime}bC$ this also implies that $y=\hat{y}\in A^{\prime}$. Furthermore, with (5),